9e^x = 107
Solve each equation for the given variable. Be certain to reject values if necessary.
I think I'm getting tripped up on the e. I think I may be able to get it if I'm shown the steps. Thank you.

Answer by fractalier(2101)

From 9e^x = 107, we first divide both sides by 9, and get
e^x = 107/9
Then take the natural logarithm of both sides
ln(e^x) = ln (107/9)
x = ln (107/9) which is about 2.476